OpenStudy (anonymous):

5x+1/2x=1/2x=10 What is x???

jhonyy9 (jhonyy9):

5x+1= - so why is there two equality 5x+1=1=10*2x 5x+1-1=20x 5x=20x 15x=0 x=0